what would the square footage of surface area of a cone shaped pond that has a radius of 100 ft and is 20 ft deep
Answer by Cromlix(4381)

Surface area of a cone = Pi* r* s
S = hypotenuse of a right angled
triangle made up of base = radius
and upright = height
20^2 + 100^2 = hypotenuse^2
Hypotenuse (s) = 101.98ft
Surface Area = Pi * 100 * 101.98
Surface Area = 32,037.96ft^2.
I have not added Pi*r^2 as the surface of the
pond is not included in the total.
Hope this helps.
:-)